40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 12 12:13:37 2020, gautam, Update, LSC, AO path attempts CARM_OLTF.pdfPRFPMI_errSigs.pdf
    Reply  Thu Feb 13 02:07:26 2020, gautam, Update, LSC, AO path transfer function measurement AOpathTF.pdfIMCinLoop.pdf
Message ID: 15208     Entry time: Wed Feb 12 12:13:37 2020     Reply to this: 15210
Author: gautam 
Type: Update 
Category: LSC 
Subject: AO path attempts 

Summary:

  1. The PRFPMI can be controlled by a mix of ALS and RF signals and circualting arm powers > 100 can be maintained for several tens of minutes at a stretch.
  2. The complete RF handoff still cannot be realized - I need to study the AO path crossover more carefully to understand what exactly is wrong and what needs to be done to rectify the problem.

Measurements:

Over the last couple of days, I've been trying to see if I can measure the phase advance due to the AO path - however, I've been unable to do so for any combination of CM board IN1 gain and MC Servo board IN2 gain I've tried. Yesterday, I tried to understand the loop shapes I was measuring a little more, and already, I think I can't explain some features.

Attachment #1 shows the TF measured (using SR785, and the EXC_A bank of the CM board) when the CM Slow path has been engaged.

  • All CARM control in this state is digital.
  • For the CM Slow path, the digital filter includes a pole at 700 Hz, pole at 5 kHz and zero at 120 Hz (the latter two for coupled cavity pole compensation).
  • In this conditions, the arm powers are somewhat stable at ~150, but still there are fluctuations of the order of 50%.
  • The "buzzing" as the arms rapidly go in and out of resonance is no longer present though.
  • The UGF of the hybrid REFL11+ALS loop is ~200 Hz, with ~45 deg of phase margin.
  • Turning off the MC2 violin filters gives some phase back. But I don't really understand the flattening of the TF gain between ~250-500 Hz.

Attachment #2 shows error signal spectra for the in-loop PRFPMI DoFs, for a few different conditions.

  • Engaging the REFL11 digital path smooths out the excess noise in the ~30-50 Hz band, which is consistent with the fact that the arm powers stabilize somewhat.
  • However, there is some gain peaking around ~400 Hz.
  • This is in turn imprinted on the vertex DoFs, making the whole system's stability marginal.

I believe that a stable crossover is hopeless under these conditions.

Next steps:

  1. Account for the measured OLTF, understand where the flattening in the few hundred Hz region is coming from.
  2. Repeat the high BW POY experiments, but with the simulated coupled cavity pole - maybe this will be a closer simulation to the PRFPMI transition.
Attachment 1: CARM_OLTF.pdf  157 kB  | Hide | Hide all
CARM_OLTF.pdf
Attachment 2: PRFPMI_errSigs.pdf  341 kB  | Hide | Hide all
PRFPMI_errSigs.pdf
ELOG V3.1.3-